# It’s Our Loss: No Privacy Amplification for Hidden State DP-SGD With Non-Convex Loss

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Privacy-preserving protocols
@@ 6. Conclusion
Our main result is that hidden state privacy amplification results for DP-SGD are not possible for all non-convex loss functions in general. To that end, we carefully construct a worst-case loss function for which the privacy analysis of DP-SGD is tight, even when the adversary can only observe the final model. However, it is currently unclear if our results extend to more commonly used (natural) non-convex loss functions in differentially private machine learning, e.g., ReLU activations with Cross Entropy loss. Specifically, it is possible that such natural non-convex loss functions satisfy additional properties that are currently unknown, which similarly result in hidden state privacy amplification. Therefore, one remaining open challenge will be to investigate whether it is possible to extract the same level of information from the final iterate of DP-SGD when used together with natural loss functions as we have been able to extract from our worst-case loss function in this work.
@@ Acknowledgements.
This work has partly been supported by a National Science Scholarship (PhD) from the Agency for Science Technology and Research, Singapore (A*STAR). We wish to thank Emiliano De Cristofaro and Jamie Hayes for providing ideas and feedback throughout the project. We also wish to thank Georgi Ganev for their feedback on early versions of the paper. Lastly, we wish to thank the reviewers for their constructive feedback, which helped to improve the paper further.

### v3

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 Workshop on Artificial Intelligence and Security ; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 Workshop on Artificial Intelligence and Security (AISec ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3689932.3694767isbn: 979-8-4007-1228-9/24/10ccs: Security and privacy Privacy-preserving protocols
@@ 6. Conclusion
Our main result is that hidden state privacy amplification results for DP-SGD are not possible for all non-convex loss functions (incl. worst-case) in general. However, it is currently unclear if our results apply specifically to more commonly used (natural) non-convex loss functions in differentially private machine learning, e.g., ReLU activations with Cross Entropy loss. For instance, it is possible that such natural non-convex loss functions satisfy additional properties that are currently unknown, which similarly result in hidden state privacy amplification. Therefore, one remaining open challenge will be to investigate whether it is possible to extract the same level of information from the final iterate of DP-SGD when used together with natural loss functions as well.
@@ Acknowledgements.
This work has partly been supported by a National Science Scholarship (PhD) from the Agency for Science Technology and Research, Singapore (A*STAR). We wish to thank Emiliano De Cristofaro and Jamie Hayes for providing ideas and feedback throughout the project. We also wish to thank Georgi Ganev for their feedback on early versions of the paper.